find vertex of x^2-6x+6
Answer by ewatrrr(24785)

 
Hi,
Parabola opening Up (coefficient of the x^2 term > 0)
y = x^2-6x+6 = (x-3)^2 -9 +6 = (x-3)^2 - 3 Vertex(3,-3)